# Unveiling Topological Structures from Language: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v3

@@ Abstract
The surge of data available on the internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 95 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001 or FA8702-25-D-B002. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2025 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work..
@@ 1 Introduction
Proliferation of the Internet has given rise to the generation of massive amounts of data. These massive amounts of data when processed can solve many crucial issues plaguing our current society. Due to this well-established notion among stake-holding institutions, the Machine Learning (ML) field has been thriving as a tool that extracts trends and solutions to non-trivial problems. However, real-world data tends to be noisy, heterogeneous, imbalanced, have missing labels, contain high-dimensionality, etc., often making the adoption of ML techniques to such datasets non-trivial. Therefore, to extract meaningful findings from data, specifically real-world data, clever techniques that extract additional features, while preserving the structure of the data need to be employed. To that end, a small niche community for Topological Data Analysis (TDA) applications in NLP has emerged. Being promised as a technique that can extract and analyze the shape/topology of data, TDA has great potential in mitigating such issues witnessed in real-world data.
@@
Geometric techniques (Delaunay triangulation, Convex hull) Uses geometrical techniques to extract characteristics of data by analyzing spacial boundaries
@@
TDA is a “collection of powerful tools that can quantify shape and structure in data” [^2] https://www.indicative.com/resource/topological-data-analysis/ and is inspired by the algebra and geometry mathematical fields. The benefits of TDA are vast, including the ability to extract additional features, typically not captured by other feature extraction techniques Uchendu et al. (2024); Michel et al. (2017). We call these features that TDA extracts - topological features. Unsurprisingly, since TDA is used to capture topological features, it has been applied to many tasks where data has distinct graphical structures Papamarkou et al. (2024); Hensel et al. (2021). These include tasks that have obvious graph-like structures, such as protein classification Dey and Mandal (2018); Lamine et al. (2023); Valeriani et al. (2024) and drug discovery Alagappan et al. (2016); to those that are not so obvious such as diabetes classification Wamil et al. (2023); Skaf and Laubenbacher (2022), image classification Horn et al. (2022); Trofimov et al. (2023), time series analysis Petri and Leitao (2020); Tymochko et al. (2021); Gholizadeh and Zadrozny (2018), and speech processing Tulchinskii et al. (2023). However, since the shape of a text is not apparent, it has not gained as much attention in Natural Language Processing (NLP) as it has done in the Computer Vision Horn et al. (2022); Trofimov et al. (2023) and Medical Singh et al. (2023); Nielson et al. (2015) domains. Still, several researchers have found ways to extract additional features using TDA that the other typical numerical representation techniques in text such as TF-IDF, Word2Vec embeddings, BERT embeddings, etc have not captured. Figure 1 shows that there has been a gradual acceleration in the number of published works in TDA for NLP applications, and we project this trend will continue in the future.
Figure 1: Number of NLP papers using TDA published each year from 2013 to May-2025
TDA aims to answer the main question - what is the true shape of a data? We survey 95 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical Karlgren et al. (2014); Port et al. (2018) and non-theoretical Zhu (2013); Doshi and Zadrozny (2018). Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@ 2 Topological Data Analysis (TDA)
Topology is defined as “the study of geometric properties and spatial relations unaffected by the continuous change of shape or size of figures,” (Oxford Dictionary). TDA is then a collection of powerful techniques that can quantify the shape and structure of data Munch (2017). While, there are alternatives to TDA (i.e., see Table 1), TDA is the only technique that can extract not only local but global features, preserve the shape and structure of the data, and is still effective even with insufficient data. Two main techniques are used to extract TDA features - Persistent Homology and Mapper.
@@ 2.1 Persistent Homology
Figure 2: Illustration of the Persistent Homology technique using different radii to find the persistent features Rieck (2020). $\epsilon$ is the ball diameter.
@@
Persistent Homology (PH) Edelsbrunner et al. (2008) is the most popular TDA technique. It uses algebraic topology methods to extract topological signatures at different spatial dimensions. This process involves representing data as a point cloud and performing deformation or perturbation processes to extract the true “shape” of data after the noise has been removed. To achieve this, PH employs Vietoris-Rips complex Munch (2017). Vietoris-Rips complex is a way to build simplicial complexes which are used to represent data in a topological space. A simplicial complex is a topological space built by putting points, lines, and higher dimensional shapes together. These formations reveal features that are holes in different dimensions, represented as betti numbers ($\beta_{d}$, $d$-dimension). Holes in the 0-dimension ($\beta_{0}$) is represented as one vertex, 1-dimension ($\beta_{1}$) is represented as an edge, 2-dimension ($\beta_{2}$) is represented as a triangle. Further, these features are called connected components, loops/tunnels, and voids, respectively.
@@
Using the method described above, data is represented as a point cloud, and circles are drawn around each point. Next, the radius of each circle is increased using a defined range of points, such that if the circles get bigger and touch, one of the points disappears and this is recorded as a $death$. Additionally, this process of perturbation in different dimensions can cause the $birth$ of a new hole which is also recorded. Thus, due to these deformations, the following TDA features can be extracted and recorded in a 3-column matrix, which consists of columns representing - the $birth$ (formation of holes), $death$ (deformation or the closing of holes), and persistence features. Persistence is defined as the length of time it took a feature to disappear or die ($death-birth$). The $death$ is recorded with the radii value at which the points overlap. Lastly, TDA features are typically visualized in a persistence diagram which is a visual representation of the 3-column matrix of TDA features. Figure 2 illustrates an example of the process of extracting TDA features using Persistent Homology. Other ways of visualizing TDA features include Persistence images Adams et al. (2017) and Barcode plots Ghrist (2008). Application-wise, PH has been used to extract novel features to complement existing NLP representations and improve various classification performances Doshi and Zadrozny (2018); Uchendu et al. (2024); Wu et al. (2022).
@@
Figure 3: Illustration of Mapper from Murugan and Robertson (2019). The filter function $f$ is a height function, which is a projection onto the y-axis. The cover of the projected space is the four intervals $U_{i}$. The Mapper graph on the right is a result of applying the rest of the Mapper algorithm and clustering each preimage in the nearest neighbor.
@@ 2.2 Mapper
Mapper is a dimension reduction clustering technique, used to visualize TDA-extracted topological structures/signatures. It was proposed by Singh et al. (2007) and has been used extensively to visualize topological structures in data to create visually pleasing figures, as well as interpret model performance through data probing Carlsson (2020). The Mapper algorithm works in 4 steps [^3] https://www.quantmetry.com/blog/topological-data-analysis-with-mapper/ (Figure 3) following Murugan and Robertson (2019)’s instructions: (1) Transform the data to a lower-dimensional space using a filter function $f$, also known as a lens. This implies projecting from one space to another. Options for filter functions include PCA Maćkiewicz and Ratajczak (1993), UMAP McInnes et al. (2018), and any other dimension-reduction algorithms; (2) Create a cover $(U_{i})_{i\in I}$ for the projected space, which is typically composed of overlapping intervals with a constant length; (3) Cluster the points in the preimage $f^{-1}(U_{i})$ into sets $C_{i,1},\ldots,C_{i,k_{i}}$ per interval $U_{i}$; (4) Create a graph where each vertex represents a cluster set, and there is an edge between two vertices if the corresponding clusters share common points. In this step, points in the same neighborhood are clustered using a defined clustering technique such as DBSCAN Ester et al. (1996) to change a cluster of several points into a node of a graph.
@@
The intrinsic nature of the Mapper algorithm makes it advantageous in preserving structure even with mapping from one dimension to another. Furthermore, the clustering techniques allow it to be used to explain model performance as the clusters, and colors have meaning that can be further explored. Finally, Mapper is more useful for exploratory data analysis, while PH is more useful for analyzing point clouds and examining the persistence of features.
@@ 3 Theoretical Approaches of TDA in NLP
Since the field of NLP is very interested in representing and analyzing texts or speech in meaningful ways, several theoretical approaches have been proposed to investigate how well these approaches align with linguistic principles. Thus to explain or confirm linguistic phenomena within the NLP paradigm, a few researchers have proposed topological approaches for probing NLP techniques. By employing TDA techniques - Persistent homology or Mapper to probe for linguistic phenomena, researchers aim to capture the topological space (both semantic and syntactic relationships) of texts, analyze and visualize topological topic evolution within texts, and extract the topological shape of words. See Table 4 in the Appendix for the theoretical approaches.
Theoretical vs. Non-Theoretical approaches. The main difference between the theoretical and non-theoretical approaches in this context is - Theoretical approaches use TDA (sometimes in combination with other numerical techniques) to understand and explain linguistic phenomena, while Non-theoretical approaches use TDA to enhance or explain model performance. However, linguistic phenomena understood from the Theoretical approaches can inform Non-theoretical approaches.

### v4

# Unveiling Topological Structures from Language: A Survey of Topological Data Analysis Applications in NLP

@@ Abstract
The surge of data available on the Internet has led to the adoption of various computational methods to analyze and extract valuable insights from this wealth of information. Among these, the field of Machine Learning (ML) has thrived by leveraging data to extract meaningful insights. However, ML techniques face notable challenges when dealing with real-world data, often due to issues of imbalance, noise, insufficient labeling, and high dimensionality. To address these limitations, some researchers advocate for the adoption of Topological Data Analysis (TDA), a statistical approach that discerningly captures the intrinsic shape of data despite noise. Despite its potential, TDA has not gained as much traction within the Natural Language Processing (NLP) domain compared to structurally distinct areas like computer vision. Nevertheless, a dedicated community of researchers has been exploring the application of TDA in NLP, yielding 100 papers we comprehensively survey in this paper. Our findings categorize these efforts into theoretical and non-theoretical approaches. Theoretical approaches aim to explain linguistic phenomena from a topological viewpoint, while non-theoretical approaches merge TDA with ML features, utilizing diverse numerical representation techniques. We conclude by exploring the challenges and unresolved questions that persist in this niche field. Resources and a list of papers on this topic can be found at: https://github.com/AdaUchendu/AwesomeTDA4NLP  [^1] DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Department of the Air Force under Air Force Contract No. FA8702-15-D-0001 or FA8702-25-D-B002. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Air Force. © 2025 Massachusetts Institute of Technology. Delivered to the U.S. Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014). Notwithstanding any copyright notice, U.S. Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above. Use of this work other than as specifically authorized by the U.S. Government may violate any copyrights that exist in this work..
@@ 1 Introduction
Proliferation of the Internet has given rise to the generation of massive amounts of data. These massive amounts of data when processed can solve many crucial issues plaguing our current society. Due to this well-established notion among stake-holding institutions, the Machine Learning (ML) field has been thriving as a tool that extracts trends and solutions to non-trivial problems. However, real-world data tends to be noisy, heterogeneous, imbalanced, have missing labels, contain high-dimensionality, etc., often making the adoption of ML techniques to such datasets non-trivial. Therefore, to extract meaningful findings from data, specifically real-world data, clever techniques that extract additional features, while preserving the structure of the data need to be employed. To that end, a small niche community for Topological Data Analysis (TDA) applications in NLP has emerged. Being promised as a technique that can extract and analyze the shape/topology of data, TDA has great potential in mitigating such issues witnessed in real-world data. Thus by applying TDA to NLP, we obtain “topological structures from language,” which refers not to intrinsic properties of raw text itself, but to the structures that emerge when linguistic data is mapped into high-dimensional embedding spaces. These induced topologies capture relationships among words, sentences, or documents based on their learned representations, rather than any inherent topological features of the text. In addition, the phrase also helps us distinguish the application of TDA and other topological approaches applied to other fields, such as computer vision from NLP.
Figure 1: Number of NLP papers using TDA published each year from 2012 to July-2025
TDA is a “collection of powerful tools that can quantify shape and structure in data” [^2] https://www.indicative.com/resource/topological-data-analysis/ and is inspired by the algebraic topology and geometry mathematical fields. The benefits of TDA are vast, including the ability to extract additional features, typically not captured by other feature extraction techniques (Uchendu et al., 2024; Michel et al., 2017; Papamarkou et al., 2024). These features are known as topological features. Unsurprisingly, since TDA is used to capture topological features, it has been applied to many tasks where data has distinct graphical structures (Papamarkou et al., 2024; Hensel et al., 2021). These include tasks that have obvious graph-like structures, such as protein classification (Dey & Mandal, 2018; Lamine et al., 2023; Valeriani et al., 2024) and drug discovery (Alagappan et al., 2016); to those that are not so obvious such as diabetes classification (Wamil et al., 2023; Skaf & Laubenbacher, 2022), image classification (Horn et al., 2022; Trofimov et al., 2023), and time series analysis (Petri & Leitao, 2020; Tymochko et al., 2021; Gholizadeh & Zadrozny, 2018). However, since the shape of a text is not apparent, it has not gained as much attention in Natural Language Processing (NLP) as it has done in the Computer Vision (Horn et al., 2022; Trofimov et al., 2023) and Medical (Singh et al., 2023; Nielson et al., 2015) domains. Still, several researchers have found ways to extract unique features using TDA, in which other typical numerical representation techniques in text, such as TF-IDF, Word2Vec embeddings, BERT embeddings, etc., cannot extract.
Furthermore, while these standard numerical representations can perform well in classifying many benchmark datasets, they are not always robust to more realistic datasets that tend to have harder constraints, such as heterogeneity of labels, imbalanced, noisy, missing data, unlabeled data, etc. Since one of the benefits of TDA is to find the features after deformation techniques have been applied to the data, it tends to be more robust to these harder constraints. Therefore, for this reason, the small niche TDA applications in NLP community has explored several ways to extract additional features for classification, interpreting/explaining model performance, and explaining linguistic phenomena within text or speech. The first application of TDA in NLP was published in 2012 (Wagner et al., 2012), and since then there have been over 90 papers applying TDA in NLP. See Figure 1 for the gradual acceleration in the number of published works in TDA applications on NLP tasks, and we project that this trend will continue in the future. All these papers will be discussed in the survey.
TDA aims to answer the main question - what is the true shape of a data? We survey 100 papers that have attempted to find an answer through various approaches. We categorize all observed approaches that incorporate TDA in NLP into two categories, namely theoretical (Karlgren et al., 2014; Port et al., 2018) and non-theoretical (Zhu, 2013; Doshi & Zadrozny, 2018) approaches. Theoretical approaches involve using TDA to explain linguistic phenomena by probing the topological space, shape, and evolution of topics. On the other hand, Non-theoretical approaches mainly discuss how to effectively apply existing numerical representation techniques in NLP to extract novel topological features with TDA. We will first discuss the principles behind TDA and the two main techniques employed for TDA feature extraction: Persistent Homology and Mapper.
@@
Geometric techniques (Delaunay triangulation, Convex hull) Uses geometrical techniques to extract characteristics of data by analyzing spatial boundaries
@@ 2 Topological Data Analysis (TDA)
Topology is defined as “the study of geometric properties and spatial relations unaffected by the continuous change of shape or size of figures,” (Oxford Dictionary). TDA is then a collection of powerful techniques that can quantify the shape and structure of data (Munch, 2017). While there are alternatives to TDA (i.e., Table 1), TDA is the only technique that can extract not only local but global features, preserving the shape and structure of the data, and is also robust to insufficient data. Two main techniques are used to extract TDA features: Persistent Homology and Mapper.
@@ 2.1 Persistent Homology
Figure 2: Illustration of the Persistent Homology technique using different radii to find the persistent features (Rieck, 2020). $\epsilon$ is the ball diameter.
@@
Persistent Homology (PH) (Edelsbrunner et al., 2008) is the most popular TDA technique. It uses algebraic topology methods to extract topological signatures at different spatial dimensions. This process involves representing data as a point cloud and performing deformation or perturbation processes to extract the true “shape” of data after the noise has been removed. To achieve this, PH employs Vietoris-Rips complex (Munch, 2017). Vietoris-Rips complex is a way to build simplicial complexes which are used to represent data in a topological space. A simplicial complex is a topological space built by putting points, lines, and higher dimensional shapes together. These formations reveal features that are holes in different dimensions, represented as betti numbers ($\beta_{d}$, $d$-dimension). Holes in the 0-dimension ($\beta_{0}$) is represented as one vertex, 1-dimension ($\beta_{1}$) is represented as an edge, 2-dimension ($\beta_{2}$) is represented as a triangle. Further, these features are called connected components, loops/tunnels, and voids, respectively.
@@
Using the method described above, data is represented as a point cloud, and circles are drawn around each point. Next, the radius of each circle is increased using a defined range of points, such that if the circles get bigger and touch, one of the points disappears and this is recorded as a deathdeath. Additionally, this process of perturbation in different dimensions can cause the birthbirth of a new hole which is also recorded. Thus, due to these deformations, the following TDA features can be extracted and recorded in a 3-column matrix, which consists of columns representing - the birthbirth (formation of holes), deathdeath (deformation or the closing of holes), and persistence features. Persistence is defined as the length of time it took a feature to disappear or die (deathbirthdeath-birth). The deathdeath is recorded with the radii value at which the points overlap. Lastly, TDA features are typically visualized in a persistence diagram which is a visual representation of the 3-column matrix of TDA features. Figure 2 illustrates an example of the process of extracting TDA features using Persistent Homology. Other ways of visualizing TDA features include persistence images (Adams et al., 2017) and barcode plots (Ghrist, 2008). In terms of application, PH has been used to extract novel features to complement existing NLP representations and improve various classification performances (Doshi & Zadrozny, 2018; Uchendu et al., 2024; Wu et al., 2022).
Persistent Homology. This is a TDA technique that studies the deformation of "holes" in different dimensions. Using PH, we can track when features appear and disappear and visualize these features, usually in a persistence diagram. This process allows us to find the true structure of data, typically devoid of noise.
@@
Figure 3: Illustration of Mapper from Murugan & Robertson (2019). The filter function $f$ is a height function, which is a projection onto the y-axis. The cover of the projected space is the four intervals $U_{i}$. The Mapper graph on the right is a result of applying the rest of the Mapper algorithm and clustering each preimage in the nearest neighbor.
@@ 2.2 Mapper
Mapper is a dimension reduction clustering technique for visualizing TDA-extracted topological structures/signatures. It was proposed by Singh et al. (2007) and has been used extensively to visualize topological structures in data to create visually pleasing figures, as well as interpret model performance through data probing (Carlsson, 2020). The Mapper algorithm works in four steps [^3] https://www.quantmetry.com/blog/topological-data-analysis-with-mapper/ (Figure 3) following Murugan & Robertson (2019)’s instructions: (1) Transform the data to a lower-dimensional space using a filter function $f$, also known as a lens. This implies projecting from one space to another. Options for filter functions include PCA (Maćkiewicz & Ratajczak, 1993), UMAP (McInnes et al., 2018), and any other dimension-reduction algorithms; (2) Create a cover $(U_{i})_{i\in I}$ for the projected space, which is typically composed of overlapping intervals with a constant length; (3) Cluster the points in the preimage $f^{-1}(U_{i})$ into sets $C_{i,1},\ldots,C_{i,k_{i}}$ per interval $U_{i}$; (4) Create a graph where each vertex represents a cluster set. There is an edge between two vertices if the corresponding clusters share common points. Points in the same neighborhood are clustered using a defined clustering technique, such as DBSCAN (Ester et al., 1996) to change a cluster of several points into a node of a graph.
@@
The intrinsic nature of the Mapper algorithm makes it advantageous in preserving structure, even with mapping from one dimension to another. Furthermore, the clustering techniques allow it to be used to explain model performance as the clusters and colors have meaning that can be further explored. Finally, Mapper is more useful for exploratory data analysis, while PH is more useful for analyzing point clouds and examining the persistence of features. In this survey, we will discuss how several researchers use Mapper to explain or enhance several phenomena in NLP tasks.
Mapper. This is a TDA technique that visualizes the graphical representation of data in order to capture the intrinsic structure. It is very useful for preserving data structure and creating visually pleasing plots which can be investigated manually to find insights.
Name Category Task TDA Technique
(Karlgren et al., 2014) Topological Space (Semantic) Identify topical density of the space Mapper
(Cavaliere et al., 2017) Topological Space (Semantic) Extracts main concepts from text Persistent Homology
(Wagner et al., 2012) Topological Space (Semantic) Analyzes document similarities Persistent Homology
(Port et al., 2018) Topological Space (Syntactic) Analyzes syntactic parameters of different language families Persistent Homology
(Port et al., 2022) Topological Space (Syntactic) Explains linguistic structures with homoplasy phenomena Persistent Homology
(Sami & Farrahi, 2017) Topology of Topic Evolution Topic evolution within documents Persistent Homology
(Draganov & Skiena, 2024) Topological “Shape” of Words Investigates the “shape” of language phylogenies in the Indo-European language family Persistent Homology
(Fitz, 2022) Topological “Shape” of Words Captures grammatical structure expressed by corpus using word manifold Persistent Homology
(Fitz et al., 2024) Topological “Shape” of Words Measures the topological complexity of Transformer-based hidden representations Persistent Homology
(Dong, 2024) Topological “Shape” of Words Investigates the topological shapes of South American languages—Nuclear-Macro-Jê (NMJ) and Quechuan families Persistent Homology
(Bouazzaoui et al., 2021) Topological “Shape” of Words Investigates the similarity between the topological shapes of the Tifinagh and Phoenician scripts Persistent Homology
Table 2: Theoretical Applications of TDA in NLP
3 Selection Criteria for papers surveyed and Taxonomy Development
3.1 Selection Criteria
In order to find all NLP papers that applied TDA, we manually searched on Google Scholar using key terms such as text mining persistent homology, checking related articles of the relevant papers, their cited papers, and different combinations of all three methods. After, obtaining over 60 papers initially, we started creating a taxonomy and categorizing the papers. Initially, we focused on TDA applications in textual data, but as we searched, we found several applications in speech, and collected such papers. Finally, we removed papers that did not apply TDA to text or human speech data. Papers removed either applied TDA to a graphical representation of reddit social networks, applied non-TDA topological techniques, or applied TDA to non-speech audio data. Using these criteria we selected only papers that fit and collected the rest following the schema.
3.2 Taxonomy Development
Based on the papers selected for the survey, we were able to categorize the applications of these papers into two approaches - theoretical and non-theoretical applications. See description below:
  • Theoretical applications of TDA in NLP: These focus on understanding, characterizing, or proving properties of language and its representations through the lens of topology. They are less about immediate performance gains and more about insight. This application aims to answer the question - “What do the shapes of embedding spaces tell us about language itself and our models of it?” Example - Analyzing embedding spaces: Using persistent homology to study whether semantic clusters, or syntactic structures, correspond to stable topological features, and finding out what that tells us about language.
  • Non-theoretical (practical) applications of TDA in NLP: These treat TDA as a tool for solving tasks, regardless of whether deeper linguistic/topological insights are obtained. The emphasis is on utility. This application aims to answer the question - “How can topological summaries directly help with applied NLP tasks?” Example - Feature engineering: Feeding persistence diagrams or topological signatures into classifiers for sentiment analysis, topic detection, or authorship attribution.
Figure 5: Illustration of the theoretical approaches researchers have employed to (1) probe texts, (2) extract TDA features, (3) use these features to explain or confirm known linguistic phenomena.
@@ 4 Theoretical Approaches of TDA in NLP
Since the field of NLP is very interested in representing and analyzing texts or speech in meaningful ways, several theoretical approaches have been proposed to investigate how well these approaches align with linguistic principles. Thus to explain or confirm linguistic phenomena within the NLP paradigm, a few researchers have proposed topological approaches for probing NLP techniques. See Figure 5 for an illustration of this pipeline. By employing TDA techniques - Persistent homology or Mapper to probe for linguistic phenomena, researchers aim to capture the topological space (both semantic and syntactic relationships) of texts, analyze and visualize the topology of topic evolution within texts, and extract the topological shape of words. See Table 2 for the theoretical approaches. In essence, these theoretical topological methods provide a conceptual bridge between linguistic theory and mathematical topology.
Theoretical vs. Non-Theoretical approaches. The main difference between the theoretical and non-theoretical approaches in this context is - Theoretical approaches use TDA (sometimes in combination with other numerical techniques) to understand and explain linguistic phenomena, while Non-theoretical approaches use TDA to enhance or explain model performance. This means that the Theoretical approaches are focused on linguistic phenomena, while Non-theoretical approaches are focused on model performance. Thus, using linguistic phenomena understood from the Theoretical approaches, can inform and enhance Non-theoretical approaches.